# An AI System for Continuous Knee Osteoarthritis Severity Grading Using Self-Supervised Anomaly Detection with Limited Data

### v1

@@ 1 Introduction
Knee Osteoarthritis (OA) is a degenerative joint disease that affects over 250 million of the world’s population [35]. Two of the most common grading systems for OA diagnosis are the Kellgren-Lawrence (KL) system [20] and the Osteoarthritis Research Society International (OARSI) atlas criteria [2, 3]. The KL system consists of five ordinal classes from grade zero to four where grade zero is healthy, grade four is severe OA and grades $\geq 2$ are the cut-off for OA diagnosis [14]. The OARSI atlas criteria diagnoses OA if the subject meets specific criteria relating to the degree of Joint Space Narrowing (JSN) and the severity of osteophytes [15]. The subjectivity of these scales have been a matter of on-going concern [1] with studies reporting wide ranges of inter-observer reliability of 0.51 to 0.89 [37] for the KL system. Low agreement between the two scales has also been reported with OA being diagnosed almost twice as often using the OARSI atlas criteria compared with the KL system [14]. The variability of OA severity grading has been attributed to the difference in the level of clinician’s experience and/or the use of subjective language such as ‘possible’ osteophytic lipping and ‘doubtful’ joint space narrowing in the guidelines for grading [13]. The challenges of the current OA grading systems suggests that ordinal classes may not be suitable for assessing OA severity and highlights the requirement for an automated system.
This work aims to overcome the weaknesses of current AI systems and models for OA grading. Firstly, the high performance of existing techniques (accuracies $\geq 90\%$) is a cause for concern given the inter-rater observer reliability between human experts can be as low as 0.51 [37], suggesting that the methods have overfit to the dataset. Secondly, several of the approaches train the model as a multi-class classification problem which do not consider the ordinal nature of the data. Thirdly, existing solutions are reliant on large datasets consisting of thousands of X-rays for training, along with ground truth OA severity labels from experts which is a tedious and costly process.
@@ 2.1 Automated Knee OA Severity Classification
Several fully-supervised Convolutional Neural Network (CNN) based approaches have been proposed to classify knee X-rays according to the KL system. Antony et al. [4] developed a technique to localise the knee joint area on the X-ray and then classify the OA severity using a CNN. The CNN simultaneously minimised a categorical cross-entropy loss and a regression mean squared error loss. Several works approved upon this performance by employing ensembles of CNNs [5] and implementing more advanced CNNs such as the ResNET [27], the HRNet [18], Siamese networks [34], Long Short-Term Memory models (LSTM) [36] and dual channel autoencoders [17]. These methods were further improved by incorporating complex data pre-processing such as image enhancement [25] and noise-reduction with Gaussian-filters, normalising with a pixel-centering method, and implementing a balanced contrast enhancement technique [32]. The majority of existing techniques optimised their performance using loss functions suitable for multi-class classification, where the distance between all grades are equidistant i.e. the distance between grade zero and grade four is equal to the distance between grade zero and grade one. Culvenor et al. [13], instead, employed an ordinal loss function to take into account the ordinal nature of KL grading system. More recently, MediAI-OA [38] has advanced beyond simple KL grade classification by automatically quantifying the degree of JSN in the medial and lateral tibiofemoral joint and detecting osteophytes in the medial distal femur, lateral distal femur, medial proximal tibia and lateral proximal tibia regions.
@@ 2.2 Anomaly Detection
AD is a field of study that aims to automatically identify data samples that differ substantially from normality. These typically involve one-class methods such as DeepSVDD [30] and PatchCore [29] that learn or extract the features of normality and detect anomalies based on their distance to the normal RS. Generative models are another class of AD models [39], particularly popular in the medical field [11, 6, 40]. At training time, they learn to reconstruct the image. At test time, they can then identify anomalies by large reconstruction errors. However, SSL based methods such as Cutpaste [23] have been dominating the field of AD in recent years, with several works leveraging SSL for medical AD tasks [10, 11]. Although the majority of AD techniques train on unlabelled data, a recent technique Dual-distribution discrepancy with self-supervised refinement (DDAD) [11] trains on both labelled and unlabelled data. They train an ensemble of reconstruction networks with the objective of modelling the normative distribution and a second ensemble of networks to model the unknown distribution. They then calculate the intra-discrepancy and inter-discrepancy of the distributions to assign anomaly scores. Finally they train a separate network via self-supervised learning to further refine the anomaly scores. They demonstrated their AD performance on chest X-rays and brain MRI.
@@ 3 Materials
Figure 1: Examples of X-rays from the training set with for each KL grade. OAKL refers to whether OA was diagnosed based on the KL scale and OAOARSI refers to whether OA was diagnosed based on the OARSI criteria.
@@ 4 Methods
4.1 Stage 1: Self-Supervised Learning
@@
The purpose of SSL is to learn a representation of healthy X-rays. There are $K$ ensembles trained on a dataset $X_{N}$ of size $N$, where $X_{N}$ is sampled from the training data, $X_{train}$ which consists of solely nominal images of healthy X-rays. In iteration $i$ of training, a data sample $x_{i}\in X_{N}$ is paired with a randomly selected data sample $x_{j}\in X_{N}$. The original FewSOME implementation relies on ImageNet pretrained weights and an ‘anchor’ for learning compact representations of the normal class where the ground truth label, $y$ is equal to zero for the duration of training. This work extends this method to use SSL where both data samples, $x_{i}$ and $x_{j}$ are input into a Stochastic Data Augmentation (SDA) module. The SDA module consists of two sets of transforms, $T_{norm}$ and $T_{anom}$. The set of transforms $T_{norm}$ are applied only to $x_{i}$ and they consist of weak, global augmentations that aim to generate additional X-rays whose representation are within the hyper-sphere of the normal RS. $T_{norm}$ consists of the identity function (i.e. no transformation), applying jitter to the image and adjusting the sharpness and brightness. The set of transforms $T_{anom}$ are applied only to $x_{j}$ and they consist of the identity function and strong augmentations, random cropping followed by resizing and Cutpaste [23]. Section 7.2 presents the results of experiments conducted to choose the optimal performing transforms for $T_{anom}$.
@@
Following the SDA, the first $l$ layers of a neural network, $\phi$, initialised with ImageNET [31] pre-trained weights are used to transform the input space, $X_{N}$ to the RS, $\phi_{l}(X_{N})\in\mathbb{R}^{c\times h\times w}$, where $c$ is the number of channels, $h$ is the height and $w$ is the width of the extracted feature maps. The later layers are removed as they are biased towards natural image classification. As the size of the training dataset is small, both $\phi_{l}(x_{i})$ and $\phi_{l}(x_{j})$, denoted as $\phi_{i,l}$ and $\phi_{j,l}$, are converted to patch level features, $P(\phi_{i,l})$ and $P(\phi_{j,l})$ using a sliding window of size $sw\times sw$ with $stride=1$ with resulting dimensions of $p\times c\times sw\times sw$ where $p$ is the number of patches. The patch level features also improve the sensitivity of the model to localised anomalies such as osteophytes. Feature aggregation is performed by average pooling for each patch and the resultant tensor is reshaped to create patch embedding maps, $F(P(\phi_{i,l}))$ and $F(P(\phi_{j,l}))\in\mathbb{R}^{\sqrt{p}\times\sqrt{p}\times c}$. The patch embedding map, $F(P(\phi_{i,l}))$ consists of $p$ patches at coordinates $z,k$, where each patch is denoted as $p_{i,k,z}$.
@@
This method exploits that X-rays are spatially similar and therefore, the Cosine Distance, CD is calculated only between patches $p_{i,z,k}\in F(P(\phi_{i,l}))$ and $p_{j,z,k}\in F(P(\phi_{j,l}))$ that have the same coordinates, $z,k$. The ground truth label $y_{i,z,k}\forall(z,k)$ is equal to zero if the identity transform in the SDA was applied to $x_{j}$ and $y_{i,z,k}\forall(z,k)$ is equal to one if cropping and resizing was applied. In the case of applying the CutPaste transform, the entire image is not affected and therefore, $y_{i,z,k}$ is equal to one if the receptive field of $p_{j,z,k}$ was affected by the CutPasting and zero otherwise. The CD between each patch, $p_{i,z,k}$ and $p_{j,z,k}$ is calculated as shown in equation 1 and denoted as $\hat{y}_{i,z,k}$. The difference between $\hat{y}_{i,z,k}$ and the ground truth label, $y_{i,z,k}$ are minimised using Binary Cross Entropy during training as shown in equation 2.
@@ 4.2 Stage 2: Pseudo Labelling and Denoising with CLIP
Each ensemble $K$ identifies a data instance as an anomaly if the AS is greater than $m$ times the maximum CD between all possible pairs in the training data (denoted as $CD_{max}$), where $1\leq m\leq\infty$. Data instances that have been voted as an anomaly across all ensembles are then assumed to be severe OA cases. The model can be retrained on these cases to improve the sensitivity to OA specific anomalies with the newly assigned pseudo label of $y=1$. However, knee X-rays often contain anomalies unrelated to the OA grade, such as the presence of screws and metal in the X-ray due to knee replacement implants [16]. Using the pre-trained CLIP [28] model with ViT-B/32 Transformer architecture, the similarity between each image and a single compositional statement,‘there is a screw present in the image’, was obtained. Based on experiments on validation data, an assumption is made that approximately 5% of the labelled training data is affected by the presence of metal in the dataset. Therefore, images with a similarity score greater than the similarity score of the $95^{th}$ percentile of the training data are considered to be ‘false’ anomalies and they are removed from the pseudo labels, resulting in a dataset $X_{d}$ of denoised pseudo labels. This stage is conducted only during training. No X-rays are removed during the dataset at test time.
@@ 4.3 Stage 3: Dual Centre Representation Learning with FewSOME
Figure 4: Stage 3 uses DC to learn two representation spaces, one for normal instances, $RS_{norm}$ and another for anomalies $RS_{OA}$ (or $RS_{sev}$). The figure also shows a hypothetical representation space demonstrating where the pseudo labels are obtained for DCRL-FSsev and DCRL-FSOA.
@@
Given there are now anomalous psuedo labels available for training, this section proposes Dual Centre Representation Learning (DCRL) with FewSOME. The goal of DCRL is to simultaneously learn two separate RSs for normality and anomalies. DCRL has the same methodology as Stage 1 SSL except the SDA is switched off and the value of $y$ is determined by where $x_{j}$ is sampled from. For example, in iteration $i$ of training, $x_{i}\in X_{N}$ is paired with a randomly selected data sample $x_{j}$ where $y_{i}=0$ if $x_{j}\in X_{N}$ and $y_{i}=1$ if $x_{j}\in X_{d}$. The objective of the training is to transform $x_{i}$ and $x_{j}$ to the RS, calculate the CD between the two instances and compare this to the ground truth label $y_{i}$.
@@
It was found that the model learns a more robust representation of healthy X-rays when training with pseudo labels of varying OA severity, while it is less robust when exposed to only pseudo labels of severe OA cases (cases with high AS scores output from the model). However, training the model on pseudo labels of more severe OA cases (cases with high AS) results in a highly accurate severe OA detector but a less accurate OA detector. To combat this trade-off, two models are trained, DCRL-FSsev and DCRL-FSOA.
@@
DCRL-FSsev is used to learn a RS for healthy cases, $RS_{norm}$ and another for severe OA cases, $RS_{sev}$. DCRL-FSOA learns a RS for normal OA cases, $RS_{norm}$ and another for all other varying severity OA cases, $RS_{OA}$. The margin, $m$ controls the level of severity of the pseudo labels. Therefore $m$ is set higher when training DCRL-FSsev and lower when training DCRL-FSOA. DCRL-FSOA can be trained iteratively, meaning it can be used to predict additional pseudo labels, which are denoised using the method of Stage 2 and then retrained on the newly assigned pseudo labels.
@@
The models are combined to develop the continuous automated grading system, DCRL-FScomb. The CD to the centre $C_{norm}$ of $RS_{norm}$ (obtained from training DCRL-FSsev) is denoted as $d_{1}$ and the CD to $C_{sev}$ of $RS_{sev}$ is denoted as $d_{2}$. The anomaly scoring function, $S_{DCRL_{sev}}$ is then equal to the difference between the distances $d_{1}$ and $d_{2}$ as can be seen from equations 5 to 7. Similarly, $S_{DCRL_{OA}}$ is calculated as in equation 8, where $d_{3}$ is the CD to the centre $C_{norm}$ of $RS_{norm}$ (obtained from training DCRL-FSOA) and $d_{4}$ is the CD to the centre $C_{OA}$ of $RS_{OA}$. The DCRL-FSsev behaves as a severe OA detector, meaning if $S_{DCRL_{sev}}$ is greater than a threshold, $t$, $x_{u}$ is assigned a score of $1+S_{DCRL_{sev}}(x_{u})$, as $0\leq S_{DCRL_{OA}}(x_{u})\leq 1$. Otherwise it is equal to $S_{DCRL_{OA}}(x_{u})$, as shown in equation 9.
@@
$S_{DCRL_{sev}}=||d_{1}-d_{2}||$ (7)
$S_{DCRL_{OA}}=||d_{3}-d_{4}||$ (8)

### v2

# An AI System for Continuous Knee Osteoarthritis Severity Grading: An Anomaly Detection Inspired Approach with Few Labels

@@ 1 Introduction
Knee Osteoarthritis (OA) is a degenerative joint disease that affects over 250 million of the world’s population [42]. Two of the most common grading systems for OA diagnosis are the Kellgren-Lawrence (KL) system [23] and the Osteoarthritis Research Society International (OARSI) atlas criteria [2, 3]. The KL system consists of five ordinal classes from grade zero to four where grade zero is healthy, grade four is severe OA and grades $\geq 2$ are the cut-off for OA diagnosis [16]. The OARSI atlas criteria diagnoses OA if the subject meets specific criteria relating to the degree of Joint Space Narrowing (JSN) and the severity of osteophytes [17]. The subjectivity of these scales has been a matter of on-going concern [1] with studies reporting wide ranges of inter-observer reliability of 0.51 to 0.89 [45] for the KL system. Low agreement between the two scales has also been observed, with [16] reporting that the cut-off for defining knee OA using the two systems should not be considered comparable. The variability of OA severity grading has been attributed to the difference in the level of clinician’s experience and/or the use of subjective language such as ‘possible’ osteophytic lipping and ‘doubtful’ joint space narrowing in the guidelines for grading [15]. The challenges of the current OA grading systems suggests that ordinal classes may not be suitable for assessing OA severity and highlights the requirement for an automated system.
This work aims to overcome the weaknesses of current AI systems and models for OA grading. Firstly, the high performance of existing techniques (accuracies $\geq 90\%$) [29, 37] is a cause for concern given the inter-rater reliability between human experts can be as low as 0.51 [45], suggesting that the methods have overfit to the dataset. Secondly, several of the approaches train the model as a multi-class classification problem which do not consider the ordinal nature of the data [43, 19, 20]. Thirdly, existing solutions are reliant on large datasets consisting of thousands of X-rays for training, along with ground truth OA severity labels from experts which is a tedious and costly process [5, 41, 46].
@@ 2.1 Automated Knee OA Severity Classification
Several fully-supervised Convolutional Neural Network (CNN) based approaches have been proposed to classify knee X-rays according to the KL system. Antony et al. [4] developed a technique to localise the knee joint area on the X-ray and then classify the OA severity using a CNN. The CNN simultaneously minimised a categorical cross-entropy loss and a regression mean squared error loss. Several works improved upon this performance by employing ensembles of CNNs [5] and implementing more advanced CNNs such as the ResNET [31], ConvNeXt [22], EfficientNetV2-S [39] and the HRNet coupled with an attention mechanism to filter out the counterproductive features [20]. Siamese networks [41] have also been employed by first localising the knee joint area and training an ensemble of classifiers for the lateral and medial components separately. Wahyuningrum et al. [43] manually preprocesses the data by cropping the X-rays to the knee joint area. They then use a CNN to extract features, followed by a Long Short-Term Memory model (LSTM) to grade the OA severity. Farooq et al. [19] makes use of both labelled and unlabelled data by first training a dual-channel adversarial autoencoder in an unsupervised manner for the task of reconstruction. They then incorporate two supervised classification networks for KL-grade and leg-side classification achieving an average F1 score of 0.76. Bose et al. [10] demonstrated the importance of precise feature selection for knee OA diagnosis. They use a CNN to extract features, followed by feature selection techniques such as Particle Swarm Optimization (PSO) and Genetic Bee Colony (GBC) to identify the most relevant features. These methods were further improved by incorporating complex data pre-processing such as image enhancement [29], resulting in an average accuracy of 91% and noise-reduction with Gaussian-filters, normalising with a pixel-centring method, and implementing a balanced contrast enhancement technique [37]. The majority of existing techniques optimised their performance using loss functions suitable for multi-class classification, where the distance between all grades are equidistant i.e. the distance between grade zero and grade four is equal to the distance between grade zero and grade one. Culvenor et al. [15], instead, employed an ordinal loss function to take into account the ordinal nature of KL grading system.
MediAI-OA [46] has advanced beyond simple KL grade classification by automatically quantifying the degree of JSN in the medial and lateral tibiofemoral joint and detecting osteophytes in the medial distal femur, lateral distal femur, medial proximal tibia and lateral proximal tibia regions. They employ a four step approach. Firstly, they train a RetinaNet [33] on bounding box annotations to detect the knee joint and regions associated with osteophytes. Secondly, they segment the joint space region by training the HRNet [44] to detect the most medial and lateral corner points of the distal femur and proximal tibia. They then train an image classification network, NASNet [49] to determine the presence of ostephytes. MediAI-OA also trained a NASNet to classify the KL grade.
@@ 2.2 Anomaly Detection
AD is a field of study that aims to automatically identify data samples that differ substantially from normality. AD algorithms train on a dataset, $X_{train}$ which consists of solely normal images. A network $\phi$ learns to extract the features of normality from the training set through various methods such as one-class, generative or Self-Supervised techniques. Given, a test instance $x_{t}$, the trained network $\phi$ can identify if it belongs to the normal or anomalous class. One-class methods such as DeepSVDD [35] and PatchCore [34] use $\phi$ to learn a RS for normality. They can then identify anomalies based on their distance to normal RS, where large distances indicate that the case is an anomaly and small distances indicate that the case is normal. Generative models are another class of AD models [47], particularly popular in the medical field [12, 6, 48]. At training time, they learn to reconstruct the image. At test time, they can then identify anomalies by large reconstruction errors. However, SSL based methods such as Cutpaste [27] have been dominating the field of AD in recent years, with several works leveraging SSL for medical AD tasks [11, 12]. Although the majority of AD techniques train on unlabelled data, a recent technique Dual-distribution discrepancy with self-supervised refinement (DDAD) [12] trains on both labelled and unlabelled data. They train an ensemble of reconstruction networks with the objective of modelling the normative distribution and a second ensemble of networks to model the unknown distribution. They then calculate the intra-discrepancy and inter-discrepancy of the distributions to assign anomaly scores. Finally they train a separate network via self-supervised learning to further refine the anomaly scores. They demonstrated their AD performance on chest X-rays and brain MRI.
@@ 3 Materials
Figure 1: Examples of X-rays from the training set for each KL grade. OAKL refers to whether OA was diagnosed based on the KL scale and OAOARSI refers to whether OA was diagnosed based on the OARSI criteria.
@@ 4 Methods
4.1 Stage 1: Self-Supervised Learning with FewSOME (SS-FewSOME)
@@
The purpose of Stage 1, Self-Supervised FewSOME (SS-FewSOME) is to learn a representation of healthy X-rays. There are $K$ ensembles trained on a dataset $X_{N}$ of size $N$, where $X_{N}$ is sampled from the training data, $X_{train}$ which consists of solely nominal images of healthy X-rays. In iteration $i$ of training, a data sample $x_{i}\in X_{N}$ is paired with a randomly selected data sample $x_{j}\in X_{N}$. The original FewSOME implementation relies on ImageNet pretrained weights and an ‘anchor’ for learning compact representations of the normal class where the ground truth label, $y$ is equal to zero for the duration of training. This work extends this method to use SSL where both data samples, $x_{i}$ and $x_{j}$ are input into a Stochastic Data Augmentation (SDA) module. The SDA module consists of two sets of transforms, $T_{norm}$ and $T_{anom}$. The set of transforms $T_{norm}$ are applied only to $x_{i}$ and they consist of weak, global augmentations that aim to generate additional X-rays whose representation are within the hyper-sphere of the normal RS. $T_{norm}$ consists of the identity function (i.e. no transformation), applying jitter to the image and adjusting the sharpness and brightness. The set of transforms $T_{anom}$ are applied only to $x_{j}$ and they consist of the identity function and strong augmentations, random cropping followed by resizing and Cutpaste [27]. Section 8.2 presents the results of experiments conducted to choose the optimal performing transforms for $T_{anom}$.
@@
Following the SDA, the first $l$ layers of a neural network, $\phi$, initialised with ImageNET [36] pre-trained weights are used to transform the input space, $X_{N}$ to the RS, $\phi_{l}(X_{N})\in\mathbb{R}^{c\times h\times w}$, where $c$ is the number of channels, $h$ is the height and $w$ is the width of the extracted feature maps. The later layers are removed as they are biased towards natural image classification. As the size of the training dataset is small, both $\phi_{l}(x_{i})$ and $\phi_{l}(x_{j})$, denoted as $\phi_{i,l}$ and $\phi_{j,l}$, are converted to patch level features, $P(\phi_{i,l})$ and $P(\phi_{j,l})$ using a sliding window of size $sw\times sw$ with $stride=1$ with resulting dimensions of $p\times c\times sw\times sw$ where $p$ is the number of patches. The patch level features also improve the sensitivity of the model to localised anomalies such as osteophytes. Feature aggregation is performed by average pooling for each patch and the resultant tensor is reshaped to create patch embedding maps, $F(P(\phi_{i,l}))$ and $F(P(\phi_{j,l}))\in\mathbb{R}^{\sqrt{p}\times\sqrt{p}\times c}$. The patch embedding map, $F(P(\phi_{i,l}))$ consists of $p$ patches at coordinates $z,k$, where each patch is denoted as $p_{i,z,k}$.
@@
This method exploits that X-rays are spatially similar and therefore, the Cosine Distance, CD is calculated only between patches $p_{i,z,k}\in F(P(\phi_{i,l}))$ and $p_{j,z,k}\in F(P(\phi_{j,l}))$ that have the same coordinates, $z,k$. The ground truth label $y_{i,z,k}\forall(z,k)$ is equal to zero if the identity transform in the SDA was applied to $x_{j}$ and $y_{i,z,k}\forall(z,k)$ is equal to one if cropping and resizing was applied. In the case of applying the CutPaste transform, the entire image is not affected and therefore, $y_{i,z,k}$ is equal to one if the receptive field of $p_{j,z,k}$ was affected by the CutPasting and zero otherwise. The CD between each patch, $p_{i,z,k}$ and $p_{j,z,k}$ is calculated as shown in equation 1 and denoted as $\hat{y}_{i,z,k}$. The difference between $\hat{y}_{i,z,k}$ and the ground truth label, $y_{i,z,k}$ are minimised using Binary Cross Entropy during training as shown in equation 2. Stage 1 is depicted in figure 2.
@@ 4.2 Stage 2: Pseudo Labelling and Denoising with CLIP
Each ensemble $K$ identifies a data instance as an anomaly if the AS is greater than $m$ times the maximum CD between all possible pairs in the training data (denoted as $CD_{max}$), where $1\leq m\leq\infty$. Data instances that have been voted as an anomaly across all ensembles are then assumed to be severe OA cases. The model can be retrained on these cases to improve the sensitivity to OA specific anomalies with the newly assigned pseudo label of $y=1$. However, knee X-rays often contain anomalies unrelated to the OA grade, such as the presence of screws and metal in the X-ray due to knee replacement implants [18]. Using the pre-trained CLIP [32] model with ViT-B/32 Transformer architecture, the similarity between each image and a single compositional statement,‘there is a screw present in the image’, was obtained. This statement was chosen based on the analysis in section 8.1.1. The cosine similarity was calculated between the compositional statement and all X-rays in the validation data. To identify a threshold, $t_{c}$ for detecting screws in X-rays, the lowest similarity score of the subset in the validation set that contained a screw was found, resulting in $t_{c}=23.05$. In Stage 2, X-rays with CLIP similarity scores $>t_{c}$ were identified as ‘false’ anomalies. These are removed from the pseudo labels, resulting in a dataset $X_{d}$ of denoised pseudo labels. This stage is conducted only during training. No X-rays are removed from the dataset at test time.
@@ 4.3 Stage 3: Dual Centre Representation Learning with FewSOME
Figure 4: Stage 3 uses DCRL to learn two representation spaces, one for normal instances, $RS_{norm}$ and another for anomalies $RS_{OA}$ (or $RS_{sev}$). The figure also shows a hypothetical representation space demonstrating where the pseudo labels are obtained for DCRL-FSsev and DCRL-FSOA.
@@
Given there are now anomalous pseudo labels available for training, this section proposes Dual Centre Representation Learning (DCRL) with FewSOME. The goal of DCRL is to simultaneously learn two separate RSs for normality and anomalies. DCRL has the same methodology as Stage 1 except the SDA is switched off and the value of $y$ is determined by where $x_{j}$ is sampled from. For example, in iteration $i$ of training, $x_{i}\in X_{N}$ is paired with a randomly selected data sample $x_{j}$ where $y_{i}=0$ if $x_{j}\in X_{N}$ and $y_{i}=1$ if $x_{j}\in X_{d}$. The objective of the training is to transform $x_{i}$ and $x_{j}$ to the RS, calculate the CD between the two instances and compare this to the ground truth label $y_{i}$.
@@
It was found that the model learns a more robust representation of healthy X-rays when training with pseudo labels of varying OA severity, while it is less robust when exposed to only pseudo labels of severe OA cases (cases with high AS scores output from the model). However, training the model on pseudo labels of more severe OA cases (cases with high AS) results in a highly accurate severe OA detector but a less accurate OA detector. This trade-off can be combatted by training two models, DCRL-FSsev and DCRL-FSOA.
@@
DCRL-FSOA is the continuous disease severity grading system that learns a RS for healthy cases, $RS_{norm}$ and another for all other varying severity OA cases, $RS_{OA}$. DCRL-FSsev is used to learn a RS for healthy cases, $RS_{norm}$ and another for severe OA cases, $RS_{sev}$ and behaves as a severe OA detector. The margin, $m$, described in section 4.2, controls the level of severity of the pseudo labels. Therefore $m$ is set higher when training DCRL-FSsev and lower when training DCRL-FSOA. DCRL-FSOA can be trained iteratively, meaning it can be used to predict additional pseudo labels, which are denoised using the method of Stage 2 and then retrained on the newly assigned pseudo labels.
@@
The models can then be combined to improve the OA and severe OA detection accuracy. This model is denoted as DCRL-FScomb. The CD to the centre $C_{norm}$ of $RS_{norm}$ (obtained from training DCRL-FSsev) is denoted as $d_{1}$ and the CD to $C_{sev}$ of $RS_{sev}$ is denoted as $d_{2}$. The anomaly scoring function, $S_{DCRL_{sev}}$ is then equal to the difference between the distances $d_{1}$ and $d_{2}$ as can be seen from equations 5 to 7. Similarly, $S_{DCRL_{OA}}$ is calculated as in equation 8, where $d_{3}$ is the CD to the centre $C_{norm}$ of $RS_{norm}$ (obtained from training DCRL-FSOA) and $d_{4}$ is the CD to the centre $C_{OA}$ of $RS_{OA}$. Given the range of possible output scores for $d1$, $d2$, $d3$ and $d4$ is $[0,2]$, the range of possible output values from $S_{DCRL_{OA}}$ and $S_{DCRL_{sev}}$ is $[-2,2]$. These scores are normalised to be between zero and one by adding the minimum value of the range, two and adding the inter-range value of four. The DCRL-FSsev behaves as a severe OA detector, meaning if $S_{DCRL_{sev}}$ is greater than a threshold, $t$, $x_{u}$ is assigned a score of $1+S_{DCRL_{sev}}(x_{u})$, as $0\leq S_{DCRL_{OA}}(x_{u})\leq 1$. Otherwise it is equal to $S_{DCRL_{OA}}(x_{u})$, as shown in equation 9.
@@
$S_{DCRL_{sev}}=\frac{(d_{1}-d_{2})+2}{4}$ (7)
$S_{DCRL_{\mathrm{OA}}}=\frac{(d_{3}-d_{4})+2}{4}$ (8)
@@
The output of $S_{DCRL_{OA}}$ is continuous with values in the range $[0,1]$ after normalisation. The AS of the combined model, DCRL-FScomb, designed to improve OA and severe OA detection has a discontinuity in its output values when combining $S_{DCRL_{OA}}$ and $S_{DCRL_{sev}}$ using the method in equation 9. When combining the models for $DCRL-FS_{comb}$, the scores between zero and one should be interpreted as representing severity levels from healthy to moderate, while scores $>1$ should be interpreted as representing the level of severity within severe cases. Despite the discontinuity, the ranking of scores is preserved, meaning that scores can be interpreted relative to each other. Figure 5 presents an overview of the three stage approach.
Figure 5: simplified overview of the proposed three stage approach.